The electrostatic field strength of a proton at distance r is E1. What would be the strength of the field, E2 when the distance is doubled?


1
Expert's answer
2022-02-28T13:52:15-0500

We know that

Electric field

"E\alpha\frac{1}{r^2}"

We can written as

"\frac{E_1}{E_2}=\frac{r_2^2}{r_1^2}"

"r_2=2r_1"

Put value

"\frac{E_1}{E_2}=(\frac{2r_1}{r_1})^2=4"

"E_2=\frac{E_1}{4}"